# Evaluating the impact of racial disproportionality in Stop & Search on expressive crimes in London

### v1

@@ 2.1 Bayesian non-parametric procedure for the Neyman moment equation
$\text{CR}(p)=\left\{\begin{array}[]{cc}-2\sum_{i=1}^{n}\log(np_{i}),&\lambda=0\\ &\\ 2n\sum_{i=1}^{n}p_{i}\log(np_{i}),&\lambda=-1\\ &\\ \frac{2}{\lambda(1+\lambda)}\sum_{i=1}^{n}\left[(np_{i})^{-\lambda}-1\right],&\lambda\neq-1\text{ or }0.\\ \end{array}\right.$
Therefore, the solution in (4) is the nonparametric likelihood which seeks to reweight the sample so that it can also satisfy the moment condition (Qin and Lawless, 1994). It has been proven to possess many properties of the conventional parametric likelihood theory (Owen, 2001). A class of generalized empirical likelihood functions are studied in Imbens et al. (1998); Chernozhukov and Hong (2003); Newey and Smith (2004). In our example, we want to place a prior on the treatment effect parameter, $\beta$, directly, and update this prior in light of the data observed. Therefore, we can replace $f\left(z_{i}\left|\beta\right.\right)$ with the profile likelihood, $p_{i}$, and then the posterior distribution for $\beta$ becomes
@@
$p_{i}=\left\{\begin{array}[]{cc}\frac{1}{n}\left[1+s+t\psi(z_{i};\beta)\right]^{-1/(1+\lambda)},&\lambda\neq-1\\ &\\ s\exp\left[t\psi(z_{i};\beta)\right],&\lambda=-1\\ \end{array}\right.$ (5)
@@
4:     Sample $\beta^{(j)}\sim\pi_{j}(\beta\left|\mathbf{z}\right.)$ using the MCMC approach, where $\pi_{j}(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\prod_{i=1}^{n}p_{i}^{(j)}$.
@@ 2.2 The role of sample splitting procedures
In Chernozhukov et al. (2018), they found that sample splitting plays a key role in reducing the bias when estimating $\beta$. The remainder term in the Taylor expansion involves the product between the error term in the PS model and the bias using ML method to estimate $\mu(\cdot)$. In some cases, this remainder term might not vanish when $n\to\infty$ as the two terms are correlated. In conventional semi-parametric analysis, we can impose Donsker conditions to restrict the class of functions that contains the estimator of $\mu(\cdot)$ so that the remainder term will be negligible. But when using ML methods where $p$ is modelled as $n$ increases, Donsker conditions are inappropriate. In Figure 1, we replicate the example in Chernozhukov et al. (2018), with $\hat{\mu}(X)=\mu(X)+(Y-\mu(X))/{n^{1/3}}$ and $\hat{\pi}(X)=\pi(X)$. We use the full sample to estimate $\beta$ using (1) and Algorithm 1 with a vague prior $\mathcal{N}(0,10000)$ and $\lambda=0,-1,-1/2$, which corresponds to EL, ETEL and HD cases respectively. The first column of Figure 1 shows the results of 10,000 replicates, and the histograms of the posterior mean via the Bayesian method in Algorithm 1. In this simple example, both methods indicate some biases from the full sample approach while the Bayesian method has a slightly smaller bias. In light of this consideration, Chernozhukov et al. (2018) proposed the use of sample splitting, that is, the data are partitioned into $K$ groups. The functions $\hat{\mu}_{k}(\cdot)$ and $\hat{\pi}_{k}(\cdot)$ are estimated using all the data excluding the $k$th group. Then the double ML estimator for $\beta$ is the solution to $01/K\sum_{k=1}^{K}\mathbb{E}_{k}[\psi(Z;\beta)]=$ 0, where $\mathbb{E}_{k}(\cdot)$ is the empirical expectation over the $k$th fold of the data. This creates the independence between two terms, leading to unbiased estimation. Therefore, it is necessary to amend Algorithm 1 to incorporate the sample splitting strategy to remove the bias. In essence, we can mimic the procedure to partition the data into $K$ groups and estimate the $\mu_{k}(\cdot)$ and $\pi_{k}(\cdot)$ using all the data excluding the $k$th group and then use them to obtain the non-parametric probability $p_{i}$ in (4) with $i\in$ $k$th group only. Algorithm 2 summarizes the update algorithm to generate the posterior sample using sample splitting.
@@
8:     Sample $\beta^{(j)}\sim\pi_{j}(\beta\left|\mathbf{z}\right.)$ using the MCMC approach, where $\pi_{j}(\beta\left|\mathbf{z}\right.)\propto\pi_{0}(\beta)\times\prod_{i=1}^{n}p_{i}^{(j)}$.
@@
The second column of Figure 1 displays the results using two-fold sample splitting, i.e., $K=2$. Both the double machine learning (ML) methods and the amended Bayesian approach effectively eliminate bias from the full-sample estimation procedure. The posterior distribution obtained using Algorithm 2 exhibits similar distributional properties across 10,000 replications.
@@ 2.3 Validating the posterior uncertainty
In this section, we evaluate whether the posterior in (6) is a valid posterior inference in terms of uncertainty. As demonstrated in Chernozhukov et al. (2018), if the estimated functions of $\mu(x)$ and $\pi(x)$ converge to the true functions in probability under some regularity conditions, the estimator for $\beta$ in (2) will converge in distribution to a normal distribution, centered at $\beta$ with variance, $\mathbb{E}[(D-\pi(X))^{2}]^{-1}\mathbb{E}[(Y-\beta D-\mu(X))^{2}]$. This is the semi-parametric efficiency bound for $\beta$.
The ‘posterior’ density in Algorithm 2, $\pi(\beta|z_{1:n})$, is a non-standard posterior inference as the likelihood is replaced by the profile likelihood with plug-in estimates for $\mu(x)$ and $\pi(x)$. Despite the theoretical support of the ETEL case in Schennach (2007); Chib et al. (2018), one would want to assess if this plug-in coupled with sample-splitting strategy is valid Bayesian inference. That is, if interval $S_{\alpha}(z)$ is a designated $1-\alpha$ probability interval, a ‘posterior’ density in Algorithm 2, $\pi(\beta|z_{1:n})$, will have the property $\mathbb{P}_{\pi}\left[\beta\in S_{\alpha}(z))\right]=1-\alpha$, if $Z_{1},\ldots,Z_{n}$ are drawn from the true data generating model. Monahan and Boos (1992) proposed a notion of proper Bayesian inference by replacing the parametric likelihood with an alternative likelihood function. They stated that the ‘posterior’ density, which derives from the alternative likelihood, should follow the law of the probability deriving from the Bayes’s rule. The posterior density is defined as valid by coverage if $\mathbb{P}_{\pi}\left[\beta\in S_{\alpha}(z))\right]=1-\alpha$, if $Z_{1},\ldots,Z_{n}$ are drawn from the true data generating model. The posterior coverage set, $S_{\alpha}(z)$, resulting from a valid posterior, should achieve nominal coverage under the joint measure of $Z$ and $\theta$, that is, $P_{\pi}(\beta\in S_{\alpha}(z))$ should have expectation $1-\alpha$ for data generated under the measure $\pi_{0}(\beta)f(z|\beta)$ on $(\beta,Z)$ for every absolutely continuous prior, $\pi_{0}(\ldotp)$. To verify this property, let
@@
and if $\pi(\beta|\mathbf{z})$ is a valid posterior, then $H$ follows Uniform$(0,1)$. In practice, if we generate $\beta_{k}(k=1,\ldots,m)\sim\pi_{0}(\ldotp)$ and the data, $z_{1:n}^{(k)}$, from $f(\ldotp|\beta_{k})$, and compute the posterior according to Algorithm 2. Then we can obtain $H_{k}$ based on (7) by replacing $\beta$ with $\beta_{k}$. If the distribution of $H_{k}$ follows the uniform distribution, then the posterior distribution generated from Algorithm 2 is defined as a coverage proper posterior, yielding valid posterior inference. This evaluation method for the validity of posterior inference has also been applied to verify the correctness of Bayesian computation (see for example, Talts et al., 2018).
In light of this approach, we investigate the validity of the proposed generalized empirical likelihood approach via a simulation study using the same set-up with the sample-splitting simulation, with each $\beta_{k}$ ($k=1,\ldots,10000)$ generating from $N(1,2)$. Figure 2 contains the histograms of the simulated $H$ over 10000 simulation runs with the associated $p$-values of the Kolmogorov–Smirnov test for uniformity. For all cases, it suggests that simulation $H$ values follow uniform distribution and indicates proper posterior inference according to Monahan and Boos (1992) when $\lambda=0,-1,-1/2$. Therefore, this simulation result gives us evidence that the proposed method according to Algorithm 2 can be regarded as a valid approach for posterior inference in terms of coverage.
Figure 2: Histgrams for $H$ statistics using the Bayesian generalized empirical likelihood. $P$-values represents the Kolmogorov–Smirnov test for uniformity of $H$.
@@ 3 Simulation
  • BDR-HD: The Bayesian doubly robust high-dimension method proposed in Antonelli et al. (2022), where the propensity score and outcome are estimated via regression models with the Gaussian process prior, and then the MCMC estimate is plugged in to a doubly robust estimator. The variance is adjusted through the frequentist bootstrap so that it will achieve the nominal coverage rate.
@@
For all the Bayesian methods, we generate $5,000$ MCMC samples and $1,000$ burn-in iterations each simulation with $1,000$ simulation replicates.
@@ 3.1 Binary exposure
In this case, we consider $D$ is binary and simulate $X=(X_{1},X_{2},\ldots,X_{p})\sim\mathcal{N}_{p}\left(0,\Sigma\right)$ and $\Sigma_{ij}=1$ if $i=j$ and 0.3 otherwise, and then simulate
[…]
@@
We are interested in estimating the average treatment effect (ATE), i.e, $\beta$. In the analyses, we take $p=500$ and $n=50,200$ and place a non-informative prior, $\mathcal{N}(0,10000)$, for the ATE. Table 1 shows the performance of ETEL, EL, HD, DML across machine learning approaches (Lasso, Random Forest, and Neural Network) and BDR-HD in terms of bias, root mean squared error (RMSE), and coverage rate. For the proposed Bayesian method, it maintain low bias and RMSE with relatively high coverage rates. Specifically, EL and HD generally achieve slightly better precision than ETEL, with HD sometimes offering the lowest bias. In terms of machine learning approaches, neural networks, although effective in reducing bias, show higher RMSE and lower coverage rates compared to Lasso and Random Forest. DML method demonstrates quite similar results with the proposed Bayesian method, in line with our previous demonstration about the uncertainty quantification. BDR-HD shows similar bias as other methods, but it achieves the lowest RMSE and the highest coverage rate due to the extra bootstrap step to adjust the posterior variance. However, the running time per iteration for $n=50$ of BDR-HD is approximately five times longer than that of the proposed Bayesian method. Overall, the proposed method demonstrates a reliable balance between the statistical performances and computational intensity.
@@ 3.2 Continuous exposure
In this example, the coefficient associated with $D$ is the main regression coefficient and is the parameter of interest. We set $p=40$ and $n=40$, and in the proposed Bayesian method, we assign an informative prior, $\mathcal{N}(1,2)$, to $\beta$. Table 2 shows the results of bias, RMSE and coverage rate across all methods. Most of the methods perform well in terms of bias, ranging between $-0.01$ and $0.06$, while the RMSE varies slightly, with most values in between $0.13$ and $0.19$. In particular, Lasso-based approaches have very low bias ($0.01-0.02$), while random forest based methods show slightly higher bias. In terms of RMSE, the proposed Bayesian method displays smallest RMSEs $(0.13-0.14)$ across all ML methods. DML methods showing a slightly higher RMSE $(0.16-0.19)$, while BDR-HD has a much higher RMSE. Regarding coverage rates, EL-based methods always achieve higher rates $(90.6-94.0)$, while ETEL-based methods show a bit lower coverage $(85.8-89.9)$. BDR-HD reaches the coverage rate at 95.9, closest to the nominal level. We also notice that methods using neural network generally have lower coverage rates than methods using Lasso and random forest. Results indicate that the proposed Bayesian method, coupled with an informative prior, performs well overall in terms of balancing low bias, low RMSE, and adequate coverage rates.
@@
4 Application: Stop and Search in London, UK
4.1 Stop and Search data
In this section, we apply the proposed methodology to the Stop & Search data in London to enhance our understanding of the impact of the racial disproportionality, particularly on expressive crimes. The Stop & Search dataset consists of individual stop and search monthly records in London from January 2019 to December 2023, including date and time, street-level location, ethnicity, gender and age of the person stopped, legislation, object searched and outcome. The data have been sourced from the Metropolitan Police and City of London police force and published in the Single Online Home National Digital Team under Open Government Licence v3.0. They are available openly from https://data.police.uk/docs/method/stops-at-location/. Before publishing, the location coordinates of the stop are anonymized (detailed methods of anonymization from the publisher can be found at https://data.police.uk/about/#location-anonymisation and the age of the person stopped has been adjusted to a corresponding age group (e.g. 18-24).
Category Searched Objects Legislation
In this research, 2.6 million Stop & Search incidents have been compiled for the observation period from January 2019 to December 2023, on seven legislative grounds (Legislation in the above table) for 11 types of Searched Objects. Upon the aggregation of the data into three listed categories, drug objects related stop and search incidents took up 64% over the observation period, as compared to 12% for acquisitive objects and 24% for expressive objects. The searched subjects’ ethnic information have been collected from “officer defined ethnic group”.
4.2 Demographic data on ethnic groups
Exploratory data analysis on Stop and Search subjects’ demographics, especially the ethnic compositions, has found disproportionally distributed Stop & Search among the ethnic groups. White people took up on average 40% over time (ranged from 38% to 42%), Black people took up on average 38% during the observation period (ranged from 36% to 40%), Asian people took up on average 17% (ranged from 14% to 18%), then Others at 5% (ranged from 4% to 6%), as depicted in Figure 1.
Figure 3: Stop and Search Subjects Ethnic Groups Proportions
4.3 Outcome definition
To apply our method, we define the outcome variable as the dispropotionality index (DI) for expression crime targeting Black people in each Borough (total 33 Boroughs) and it can be calculated as:
$\text{DI}_{i}=\frac{\text{rate of black people in Stop \& Search for expression crime in Borough }i}{\text{rate of average black people in Stop \& Search for expression crime in London}}.$
This index compares the rate of Black people in the Stop & Search for expression crime in each Borough to the average rate in London:
  • If $\text{DI}=1$, it indicates no disproportionality;
  • If $\text{DI}<1$, it indicates too little Stop & Search below expectation;
  • If $\text{DI}>1$, it indicates too much Stop & Search above expectation.
Figure 4: DI for Black People in Stop and Search for Expressive Reasons
Boroughs in South-West London, such as Richmond upon Thames and Kingston upon Thames, as well as City of London, are depicted in dark red with highest DI; whilst boroughs in the east alongside the banks, such as Lewisham, Greenwich, Bexley, Newham, Barking and Dagenham, as well as the very North Enfield and far South Croydon, have presented lower DI among boroughs.
Figure 5: Monthly Stop and Search in London (2019-2023)
4.4 Results
@@
We use the percentage of black people in each borough as the treatment variable. The estimated effect is adjusted by variables such as percentage of males (gender), unemployment number and migrant rate, with a total of 31 variables. A summary of all variables, including their descriptive statistics aggregated in the borough level in the analysis, is given in Table 3. We consider the Bayesian partially linear structural equation models with an informative prior, $\mathcal{N}(0,1)$, indicating the prior belief and assumption on no racial disproportion in Stop & Search practice in London. We implement the Bayesian generalized empirical likelihood with $\lambda=0,-1,-1/2$ coupled with Lasso, random forest and neural network methods, and generate in total 10000 posterior samples in each case with 1000 burn-in iterations. Table 4 shows the results from applying the proposed method. Across all cases, the posterior mean estimates are consistently negative, ranging from -0.55 to -0.31. In addition, while it shows negligible differences across ETEL, EL and HD, the results for different machine learning methods vary, especially the width of the credible intervals. Those for Lasso are the widest, and this might be due to the high correlation among some features, leading to high variability in cross-validation when selecting the optimal tuning parameter. Random forest and neural network models have relatively narrower confidence intervals, with the narrowest under EL and HD methods among them. Overall, all methods suggest a negative relationship in terms of posterior mean, indicating that disproportionality in Stop & Search aimed at the Black community may be alleviated by taking into account the proportion of the Black population when focusing on Stop & Search against Black people for expressive crimes. That is, the higher the percentage of black residents composition in the borough, the lower the predicted value of DI of Stop & Search on Black people for expressive reasons.
Table 3: Descriptive statistics of the Stop & Search data in London
@@
Table 4: Posterior mean of the effect of for expression crime targeting black population with associated 95% credible interval.

### v2

@@ 3.3 Bayesian non-parametric procedure for the double machine learning method
[…]
@@
[…]
@@
[…]
@@ 3.4 The role of sample splitting procedures
When using ML methods to estimate the nuisance functions, Chernozhukov et al. (2018) found that sample splitting plays a key role in reducing the bias when estimating $\beta$. When showing the consistency, the remainder term in the Taylor expansion involves the product between the error term in the PS model and the bias using ML method to estimate $\mu(\cdot)$. In some cases, this remainder term might not vanish when $n\to\infty$ as the two terms are correlated. In conventional semi-parametric analysis, we can impose Donsker conditions to restrict the class of functions that contains the estimator of $\mu(\cdot)$ so that the remainder term will be negligible. But when using ML methods where $p$ is modelled as $n$ increases, Donsker conditions are inappropriate. In Figure 2, we replicate the example in Chernozhukov et al. (2018), with $\hat{\mu}(X)=\mu(X)+(Y-\mu(X))/{n^{1/3}}$ and $\hat{\pi}(X)=\pi(X)$. We use the full sample to estimate $\beta$ using (1) and Algorithm 1 with a vague prior $\mathcal{N}(0,10000)$ and $\lambda=0,-1,-1/2$, which corresponds to EL, ETEL and HD cases respectively. The first column of Figure 2 shows the results of 10,000 replicates, and the histograms of the posterior mean via the Bayesian method in Algorithm 1. In this simple example, both methods indicate some biases from the full sample approach while the Bayesian method has a slightly smaller bias. To overcome this issue, Chernozhukov et al. (2018) proposed the use of sample splitting, that is, the data are partitioned into $K$ groups. The functions $\hat{\mu}_{k}(\cdot)$ and $\hat{\pi}_{k}(\cdot)$ are estimated using all the data excluding the $k$th group. Then the double ML estimator for $\beta$ is the solution to $1/K\sum_{k=1}^{K}\mathbb{E}_{k}[\psi(Z;\beta)]=0$, where $\mathbb{E}_{k}(\cdot)$ is the empirical expectation over the $k$th fold of the data. This creates the independence between two terms, leading to unbiased estimation. Therefore, it is necessary to amend Algorithm 1 to incorporate the sample splitting strategy to remove the bias. In essence, we can mimic the procedure to partition the data into $K$ groups and estimate the $\mu_{k}(\cdot)$ and $\pi_{k}(\cdot)$ using all the data excluding the $k$th group and then use them to obtain the non-parametric probability $p_{i}$ in (4) with $i\in$ $k$th group only. Algorithm 2 summarizes the update algorithm to generate the posterior sample using sample splitting.
@@
[…]
@@
The second column of Figure 2 displays the results using two-fold sample splitting, i.e., $K=2$. Both the double ML methods and the amended Bayesian approach effectively eliminate bias from the full-sample estimation procedure. The posterior distribution obtained using Algorithm 2 exhibits similar distributional properties across 10,000 replications.
@@ 3.5 Validating the posterior uncertainty
In this section, we evaluate whether the posterior in (6) is a valid posterior inference in terms of uncertainty. As demonstrated in Chernozhukov et al. (2018), if the estimated functions of $\mu(\cdot)$ and $\pi(\cdot)$ converge to the true functions in probability under some regularity conditions, the estimator for $\beta$ in (2) will converge in distribution to a normal distribution, centered at $\beta$ with variance, $\mathbb{E}[(D-\pi(X))^{2}]^{-1}\mathbb{E}[(Y-\beta D-\mu(X))^{2}]$. This is the semi-parametric efficiency bound for $\beta$.
@@
$H=\int_{-\infty}^{\beta}\pi(\varphi|z)d\varphi,$ (7)
[…]
In light of this approach, we investigate the validity of the proposed generalized empirical likelihood approach via a simulation study using the same set-up with the sample-splitting simulation, with each $\beta_{k}$ ($k=1,\ldots,10000)$ generating from $N(1,2)$. Figure 3 contains the histograms of the simulated $H$ over 10000 simulation runs with the associated $p$-values of the Kolmogorov–Smirnov test for uniformity. For all cases, it suggests that simulation $H$ values follow uniform distribution and indicates proper posterior inference according to Monahan and Boos (1992) when $\lambda=0,-1,-1/2$. Therefore, this simulation result gives us evidence that the proposed method according to Algorithm 2 can be regarded as a valid approach for posterior inference.
Figure 3: Histograms for $H$ statistics using the Bayesian generalized empirical likelihood. $P$-values represents the Kolmogorov–Smirnov test for uniformity of $H$.
@@ 4 Simulation
  • BDR-HD: The Bayesian doubly robust high-dimension method proposed in Antonelli et al. (2022), where the propensity score and outcome are estimated via regression models with the GP prior, and then the MCMC estimate is plugged in to a doubly robust estimator. The variance is adjusted through the frequentist bootstrap so that it will achieve the nominal coverage rate.
@@
For all the Bayesian methods, we generate $5,000$ MCMC samples and $1,000$ burn-in iterations each simulation with $1,000$ simulation replicates. The code for the simulation is publicly available on our GitHub page at https://github.com/yumcgill/Bayesian-DML.
@@ 4.1 Binary exposure
In this case, we consider $D$ is binary and simulate
$X=(X_{1},X_{2},\ldots,X_{p})\sim\mathcal{N}_{p}\left(0,\Sigma\right),\Sigma_{ij}=\begin{cases}1,&\text{if }i=j,\\ 0.3,&\text{if }i\neq j,\end{cases}$
and then simulate
$D\left|X\right$.  $\sim\text{ Bernoulli}\left(\text{expit}\left(0.3X_{1}+0.2X_{2}-0.4X_{5}\right)\right)$
@@
We are interested in estimating the average treatment effect (ATE), i.e, $\beta$. In the analyses, we take $p=500$ and $n=50,200$ and place a non-informative prior, $\mathcal{N}(0,10000)$, for the ATE. Table 1 shows the performance of ETEL, EL, HD, DML across machine learning approaches (Lasso, Random Forest, and Neural Network) and BDR-HD in terms of bias, root mean squared error (RMSE), and coverage rate. For the proposed Bayesian method, it maintains low bias and RMSE with relatively high coverage rates. Specifically, EL and HD generally achieve slightly better precision than ETEL, with HD sometimes offering the lowest bias. In terms of machine learning approaches, neural networks, although effective in reducing bias, show higher RMSE and lower coverage rates compared to Lasso and Random Forest. The DML method demonstrates quite similar results with the proposed Bayesian method, in line with our previous demonstration about the uncertainty quantification. BDR-HD shows similar bias as other methods, but it achieves the lowest RMSE and the highest coverage rate due to the extra bootstrap step to adjust the posterior variance. However, the running time per iteration for $n=50$ of BDR-HD is approximately five times longer than that of the proposed Bayesian method. Overall, the proposed method demonstrates a reliable balance between the statistical performances and computational intensity.
@@ 4.2 Continuous exposure
In this example, the coefficient associated with $D$ is the main regression coefficient and is the parameter of interest. We set $p=40$ and $n=40$, and in the proposed Bayesian method, we assign a relatively informative prior, $\mathcal{N}(1,2)$, to $\beta$. Table 2 shows the results of bias, RMSE and coverage rate across all methods. Most of the methods perform well in terms of bias, ranging between 0.01-0.01 and $0.06$, while the RMSE varies slightly, with most values in between $0.13$ and $0.19$. In particular, Lasso-based approaches have very low bias (0.010.020.01-0.02), while random forest based methods show slightly higher bias. In terms of RMSE, the proposed Bayesian method displays smallest RMSEs (0.130.14)(0.13-0.14) across all ML methods. DML methods showing a slightly higher RMSE (0.160.19)(0.16-0.19), while BDR-HD has a much higher RMSE. Regarding coverage rates, EL-based methods always achieve close to nominal level rates (90.694.0)(90.6-94.0), while ETEL-based methods show a bit lower coverage (85.889.9)(85.8-89.9). BDR-HD reaches the coverage rate at 95.9, closest to the nominal level. We also notice that methods using neural network generally have lower coverage rates than methods using Lasso and random forest. Results indicate that the proposed Bayesian method with EL-based likelihood, coupled with an informative prior, performs well overall in terms of balancing low bias, low RMSE, and adequate coverage rates.
@@
5 Real data application
@@
In this section, we apply the proposed methodology to the Stop and Search data, described in Section 2 in London to enhance our understanding of the impact of the racial disproportionality, particularly on expressive crimes. To conduct an informative analysis, we assign an informative prior to the model, specifically $\mathcal{N}(0,2)$, which reflects our initial assumption that there is no effect of disproportionality in stop and search practices. In terms of confounders, we include variables such as percentage of males (gender), unemployment number and migrant rate, with a total of 31 variables. A summary of all variables, including their descriptive statistics aggregated in the borough level in the analysis, is given in Table 3. We estimate the both $\pi(\cdot)$ and $\mu(\cdot)$ using all the confounders listed in Table 3, and then apply these estimates in the Neyman-orthogonal equation via the Bayesian generalized empirical likelihood. We implement the Bayesian generalized empirical likelihood with $\lambda=0,-1,-1/2$ coupled with Lasso, random forest and neural network methods, and generate in total 10,000 posterior samples in each case with 1,000 burn-in iterations. Table 4 shows the results from applying the proposed method. Across all cases, the posterior mean estimates are consistently negative, ranging from -0.55 to -0.31. In addition, while it shows negligible differences across ETEL, EL and HD, the results for different ML methods vary, especially the width of the credible intervals. Those for Lasso are the widest, and this is due to the high correlation among some features, leading to high variability in cross-validation when selecting the optimal tuning parameter. Random forest and neural network models have relatively narrower confidence intervals, with the narrowest under EL and HD methods among them. Overall, all methods suggest a negative relationship in terms of the posterior mean, indicating that accounting for the proportion of the Black population can help alleviate disproportionality in stop and search practices targeted at the Black community. Specifically, when focusing on stop and search incidents involving Black individuals for expressive crimes, we observe that as the percentage of Black residents in a borough increases, the predicted value of the DI for these practices decreases. This finding suggests that boroughs with a higher percentage of its population composed of Black individuals will report lower levels of DI in stop and search practices.
Table 3: Descriptive statistics of the Stop and Search data in London
@@
Table 4: Posterior mean of the effect of for expressive crime involving black population with associated 95% credible interval.